(y^2-2xy)dx-(2xy-x^2)dy=0
1
Expert's answer
2020-10-27T18:58:11-0400

"({y^2} - 2xy)dx - (2xy - {x^2})dy = 0"

"- \frac{{dy}}{{dx}}( - {x^2} + 2xy) + {y^2} - 2xy = 0"

Let "y(x) = xv(x)", which gives "\frac{{dy}}{{dx}} = x\frac{{dv}}{{dx}} + v" :

"- ( - {x^2} + 2{x^2}v)\left( {x\frac{{dv}}{{dx}} + v} \right) + {x^2}{v^2} - 2{x^2}v = 0"

Simplify:

"- {x^2}\left( { - x\frac{{dv}}{{dx}} + {v^2} + v + 2x\frac{{dv}}{{dx}}v} \right) = 0"

Solve for "{\frac{{dv}}{{dx}}}":

"\frac{{dv}}{{dx}} = \frac{{ - {v^2} - v}}{{x(2v - 1)}}"

Divide both sides by "\frac{{ - {v^2} - v}}{{2v - 1}}:"

"\frac{{\frac{{dv}}{{dx}}(2v - 1)}}{{ - {v^2} - v}} = \frac{1}{x}"

Integrate both sides with respect to x:

"\int {\frac{{\frac{{dv}}{{dx}}(2v - 1)}}{{ - {v^2} - v}}} dx = \int {\frac{1}{x}} dx"

Evaluate the integrals:

"- 3\ln (v + 1) + \ln v = \ln x + C"

Substitute back:

"- 3\ln (\frac{y}{x} + 1) + \ln \frac{y}{x} = \ln x + C"